### v1

# TLSQL: Table Learning Structured Query Language

Abstract.
Table learning, which lies at the intersection of machine learning and modern database systems, has recently attracted growing attention. However, existing frameworks typically require explicit data export and extensive feature engineering, creating a high barrier for database practitioners. We present TLSQL (Table Learning Structured Query Language), a system that enables table learning directly over relational databases via SQL-like declarative specifications. TLSQL is implemented as a lightweight Python library that translates these specifications into standard SQL queries and structured learning task descriptions. The generated SQL queries are executed natively by the database engine, while the task descriptions are consumed by downstream table learning frameworks. This design allows users to focus on modeling and analysis rather than low-level data preparation and pipeline orchestration. Experiments on real-world datasets demonstrate that TLSQL effectively lowers the barrier to integrating machine learning into database-centric workflows. Our code is available at https://github.com/rllm-project/tlsql/.
Table Learning, Structured Query Language, Declarative Machine Learning, Relational Databases
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Database management system enginesccs: Information systems Query languagesccs: Computing methodologies Machine learning
1. Introduction
The rapid adoption of artificial intelligence has brought table learning (Borisov et al., 2022) to the forefront of research at the intersection of database systems and machine learning. Databases, particularly relational databases, form the backbone of modern information systems and store the vast majority of enterprise data. Concurrently, advances in relational table learning (RTL) (Zahradník et al., 2023; Robinson et al., 2024) aim to enable end-to-end machine learning directly over relational databases. Therefore, table learning is increasingly recognized as one of the most promising application scenarios for deploying machine learning in practice.
Despite these advances, making table learning accessible to non-expert users, particularly database practitioners, remains challenging. Existing table learning solutions often follow a database-decoupled workflow: data must be exported, extensively preprocessed, and fed into programmatic training pipelines. This approach not only increases system complexity but also conflicts with the SQL-centric workflow that database users have relied on for decades.
Recent developments in big data analytics demonstrate that exposing familiar SQL abstractions can significantly improve usability, as exemplified by systems like Spark SQL (Armbrust et al., 2015) and Hive (Thusoo et al., 2009). However, most SQL-like machine learning interfaces are tightly coupled to specific platforms and do not generalize to arbitrary databases, limiting their applicability in broader database-centric learning scenarios. Nevertheless, these observations suggest a key insight: enabling declarative, SQL-like specifications for table learning could substantially narrow the gap between modern databases and machine learning.
Motivated by this, we propose TLSQL (Table Learning Structured Query Language), a system that allows users to perform table learning over modern databases using SQL-like declarative syntax. TLSQL is implemented as a lightweight Python library that translates custom SQL-like statements into standard SQL queries and structured table learning task descriptions. The SQL queries are executed directly by the database engine, while the task descriptions are consumed by downstream table learning frameworks, eliminating the need for explicit data export or manual feature engineering. Consequently, users can focus on modeling and analysis rather than low-level data preparation and pipeline management.
Specifically, TLSQL enables users to declaratively define machine learning tasks over relational databases, through three primary language constructs:
  • PREDICT VALUE, which declares the prediction target and task type (e.g., classification or regression) and specifies the prediction output.
  • TRAIN WITH, which specifies the tables, columns, and filtering predicates for constructing the training dataset.
  • VALIDATE WITH, which defines validation data for model selection and hyperparameter tuning.
In summary, TLSQL establishes a practical bridge between modern databases and machine learning by introducing a SQL-like declarative paradigm for table learning task specification. By lowering the barrier to entry, TLSQL enables database users to seamlessly integrate machine learning into existing SQL-centric workflows. We believe that TLSQL offers both a novel design perspective and a practical system prototype, paving the way for future database-centric learning systems.
2. System Overview
This section presents the syntax and execution workflow of TLSQL. We first introduce the core language constructs, followed by an overview of how TLSQL statements are processed and translated into executable SQL queries and structured task descriptions.
2.1. Language Syntax
To facilitate table-centric machine learning directly over relational databases, TLSQL introduces three core statements: PREDICT, TRAIN, and VALIDATE. These statements declaratively specify the prediction task, the training dataset, and the validation dataset.
PREDICT Statement
The PREDICT statement defines the prediction target, the task type (classification or regression), and optional filtering conditions for the test set. Its syntax is as follows:
PREDICT VALUE (column_selector, TASK_TYPE)
FROM table
[WHERE conditions]
Here, column_selector can be either table.column or column; TASK_TYPE must be CLF (classification) or REG (regression). The FROM clause specifies the table containing the prediction target, and the optional WHERE clause allows row-level filtering in the test set.
TRAIN Statement
The TRAIN statement specifies the provenance of the training dataset. Its syntax is:
TRAIN WITH column_selector
FROM table1, table2, ...
[WHERE conditions]
The FROM clause may list one or more tables, separated by commas. The optional WHERE clause supports Boolean expressions across multiple tables. During execution, TLSQL groups column selectors by table and routes cross-table predicates to their corresponding tables, generating independent standard SQL queries for each table.
VALIDATE Statement
The VALIDATE statement specifies the validation dataset. Since general machine learning methods require the validation and test sets to be semantically consistent (i.e., drawn from the same sample space and label space), the syntax of VALIDATE mirrors that of PREDICT, differing only in the leading keyword:
VALIDATE WITH column_selector
FROM table
[WHERE conditions]
In this statement, the task type is omitted and implicitly inherited from the corresponding PREDICT specification.
Figure 1. TLSQL compilation pipeline illustrated using the TML1M dataset (Li et al., 2024), which comprises three interconnected tables: users, movies, and ratings.
2.2. Execution Workflow
As shown in Figure 1, TLSQL follows a three-stage processing pipeline consisting of a Lexer, a Parser, and a SQLGenerator. This modular architecture preserves syntactic compatibility with SQL while supporting multi-table specifications and robust data partitioning logic.
Lexer
The Lexer performs lexical analysis by identifying keywords (e.g., PREDICT, TRAIN, VALIDATE, FROM, WHERE), operators, and identifiers, encapsulating them into Token objects annotated with positional metadata. Invalid characters or unclosed strings trigger precise, position-aware exceptions, ensuring a structurally coherent token stream.
Parser
The Parser executes syntactic analysis by consuming the token stream and applying the TLSQL grammar. Using recursive descent, it delineates clause boundaries, resolves nested structures, validates composition, and constructs a hierarchical AST. The parser also detects and localizes syntactic errors, providing detailed diagnostics. This AST captures the logical structure of TLSQL statements and forms the basis for subsequent SQL generation.
Figure 2. End-to-end relational table learning with TLSQL.
1
SQLGenerator
The SQLGenerator traverses the AST to produce standard SQL queries and extract structured table learning task descriptions. Specifically, it:
  • Groups WITH columns by table.
  • Decomposes WHERE conditions into atomic predicates and assigns them to the corresponding tables.
  • Extracts the prediction target column and task type from the VALUE clause.
The generated SQL can be executed directly by existing database systems, while the structured task descriptions are consumed by downstream table learning frameworks. This enables a single TLSQL program to orchestrate an end-to-end workflow encompassing data extraction, task specification, and model training.
3. Demonstration
This section presents three demonstrations of TLSQL. The first demonstrates an end-to-end relational table learning workflow by integrating TLSQL with an RTL model. The second shows the basic programming interface of TLSQL. The third illustrates TLSQL’s multi-level task specification and default behaviors.
3.1. End-to-End Relational Table Learning
Figure 2 illustrates an end-to-end relational table learning workflow enabled by TLSQL, demonstrating its integration with the RTL method BRIDGE (Li et al., 2024) over a relational database. [^1] For model training and evaluation, we use the official rLLM project: https://github.com/rllm-project/rllm. In this demonstration, we continue to use the TML1M dataset, which is stored in a MySQL database. The workflow proceeds through the following five steps, as annotated in the figure:
① Table Learning Task Declaration. Users declaratively specify the learning task using TLSQL statements. In this example, the PREDICT statement defines a classification task on the Age attribute of female users. The TRAIN statement selects attributes from the users, movies, and ratings tables, while restricting the training set to male users with userID less than 3000. The VALIDATE statement further specifies a validation set consisting of male users with userID greater than 3000.
② TLSQL Compilation & SQL Execution. TLSQL converts the declarative specifications to standard SQL queries along with structured metadata describing the learning task. The SQL queries are executed directly by the MySQL engine to retrieve relevant subsets of relational table data, while the structured metadata is passed to downstream table learning frameworks.
③ Data Preprocessing & Loading. The retrieved tables are subsequently preprocessed and transformed into a relational graph representation compatible with relational table learning models. This step includes schema alignment, feature organization, and seamless data loading, and is fully automated within the TLSQL workflow.
④ Hyperparameter Configuration & Model Training. The processed data are fed into the BRIDGE model for training and evaluation. Users may explicitly configure model hyperparameters, or alternatively rely on system-provided default settings.
⑤ Visualization & Result Export. Finally, the system provides visualization and result export functionalities. Users can interactively inspect training statistics, validation performance, and prediction outcomes, enabling efficient analysis and iterative refinement of learning tasks.
3.2. Programming Interface
Figure 3 illustrates the core implementation of the TLSQL conversion system. The convert function translates a TLSQL query into executable SQL statements while simultaneously extracting structured metadata required for downstream learning tasks. Internally, the conversion follows a two-stage pipeline consisting of parsing and generation. By encapsulating the entire workflow within a single function call, TLSQL provides a concise and user-friendly programming interface.
The returned ConversionResult encapsulates the SQL queries generated for each referenced table as well as structured metadata describing the learning task. This abstraction enables users to directly submit the generated SQL queries to the underlying database system, while passing the metadata to downstream components such as feature preprocessors, data loaders, or table learning models. As a result, users can focus on high-level task formulation rather than low-level data extraction or pipeline orchestration.
1@classmethod
2def convert(tlsql: str) -> ConversionResult:
3 # Parse TLSQL to AST
4 parser = Parser(tlsql)
5 ast = parser.parse()
6
7 # Generate SQL with metadata
8 generator = SQLGenerator()
9 return generator.build(ast)
10
11if __name__ == "__main__":
12 predict_query = """
13 PREDICT VALUE(users.Age, CLF)
14 FROM users
15 WHERE users.Gender=’F’"""
16 result = convert(predict_query)
Figure 3. A minimal example of the TLSQL programming interface.
3.3. Three-Level Task Specification
This part introduces TLSQL’s three-level specification logic, as illustrated in Figure 4, where certain statements are optional. The workflow supports three levels of task specification with increasing degrees of control:
  • Level I (Minimal Specification). Users specify only the PREDICT statement. If the TRAIN statement is omitted, TLSQL automatically uses all remaining data as training data. When the VALIDATE statement is not specified, the system defaults to 5-fold cross-validation on the training data.
  • Level II (Partial Specification). Users specify both the PREDICT and TRAIN statements, while the VALIDATE statement remains optional. In this case, TLSQL applies 5-fold cross-validation by default for validation.
  • Level III (Full Specification). Users explicitly define all three statements—PREDICT, TRAIN, and VALIDATE—to gain full control over data selection, partitioning, and evaluation strategies.
This design allows users to start with minimal input and add more control when needed, making TLSQL easy to use for beginners while remaining flexible for advanced users.
Figure 4. Three-level specification in TLSQL.
1
4. Conclusion
In this work, we presented TLSQL, a declarative SQL-like interface for table learning over modern relational databases. By translating TLSQL statements into standard SQL queries and structured task descriptions consumable by downstream table learning frameworks, TLSQL eliminates the need for data export, manual preprocessing, or low-level pipeline management. We believe TLSQL provides a compelling foundation for future research in database-centric learning, offering both a conceptual framework and a practical prototype. As future work, we plan to integrate large language models to generate and refine TLSQL programs from natural language, further simplifying end-to-end table learning in real-world settings.